# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 4.3. Stochastic incompressible 2D Navier-Stokes equations
Let us prove Theorem 2.4. The deterministic incompressible 2D Navier-Stokes equation can be formulated as
@@
where $u:[0,T]\times{\mathcal{O}}\to\mathbb{R}^{2}$ denotes the velocity field of an incompressible Newtonian fluid with viscosity $\nu>0$, $p:[0,T]\times{\mathcal{O}}\to\mathbb{R}$ denotes the pressure, and $f:[0,T]\times{\mathcal{O}}\to\mathbb{R}^{2}$ denotes an external force, where ${\mathcal{O}}\subset\mathbb{R}^{2}$ is a bounded domain with sufficiently smooth boundary, see [77]. We employ Dirichlet boundary conditions. Define
$V:=\{v\in W^{1,2}_{0}({\mathcal{O}};\mathbb{R}^{2})\;\colon\;\nabla\cdot v=0\text{ a.e. in }{\mathcal{O}}\}.$
@@
Denote the extensions $A_{0}:V\to V^{\ast}$ and $F:V\times V\to V^{\ast}$ by the same respective symbols. Note that these extensions can be defined by duality for Lipschitz boundary. Consider the stochastic incompressible 2D Navier-Stokes equation with viscosity $\nu>0$ driven by degenerate additive Lévy noise
(4.12) $\mathrm{d}X_{t}=(A_{0}X_{t}+F(X_{t}))\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x,$
where $B\in L_{2}(H)$ and $D\in{\mathcal{Z}}$, and $(Z,{\mathcal{Z}},\pi)$ is a measure space as in the hypotheses. Existence and uniqueness of solutions to (2.12) have been discussed in [14, Example 2.6]. Let us verify our Hypotheses in the Gelfand triple
@@
It is well-known, see [77] or [61, Lemma 2.1] that
$\|u\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}\leq 2\|u\|_{H}^{2}\|u\|_{V}^{2},\quad u\in V.$
@@
The hemicontinuity has been proved in [14, Example 2.6].
@@
The coercivity has been proved [14, Example 2.6], with $\alpha=2$ and $\delta_{1}=2\nu$.
@@
By [61, Lemma 2.2], noting that $\langle F(u),u\rangle=0$ for $u\in V$, we get for all $u,v\in V$,
@@
$=$ $-2\langle F(u,u-v),v\rangle+2\langle F(v,u-v),v\rangle$
$=$ $-2\langle F(u-v),v\rangle$
$\leq$ $4\|u-v\|_{V}^{3/2}\|u-v\|^{1/2}_{H}\|v\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}$
$\leq$ $\nu\|u-v\|_{V}^{2}+\frac{64}{\nu^{3}}\|v\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}\|u-v\|_{H}^{2}$
as in [14, Equation (2.20)] for
$\rho(v):=\frac{64}{\nu^{3}}\|v\|_{L^{4}({\mathcal{O}};\mathbb{R}^{2})}^{4}\leq\frac{128}{\nu^{3}}\|v\|_{H}^{2}\|u\|_{V}^{2}.$
@@
$2\langle A_{0}u+F(u)-A_{0}v-F(v),u-v\rangle\leq-\nu\|u-v\|^{2}_{V}+\rho(v)\|u-v\|_{H}^{2}.$
We get that Hypothesis (B) holds with $\alpha=\beta=2$, $\delta_{2}=\nu c_{0}^{2}$ and $\rho(v)=\frac{64}{\nu^{3}}\|v\|_{L^{4}({\mathcal{O}},\mathbb{R}^{2})}^{4}$ and $C_{2}=\frac{128}{\nu^{3}}$.
@@
Hypothesis (D) holds by [14, Example 2.6].
@@
Then, Theorem 2.4 follows directly from Theorem 2.1.
@@ 4.4. Stochastic shear thickening incompressible power-law fluids
Let us prove Theorem 2.5. Consider the velocity field of a viscous and incompressible non-Newtonian fluid perturbed by Lévy noise with Dirichlet boundary conditions on a sufficiently smooth bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, with outward unit normal $n$ on $\partial{\mathcal{O}}$, $d\geq 2$. Let $p>2$, $\nu>0$, and assume that $p\geq\frac{2+d}{2}$. The case $p>2$ is called shear thickening. For $u:{\mathcal{O}}\to\mathbb{R}^{d}$, define
@@
and define the convection term $F:(W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V)\times(W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V)\to H$ as before in the case of the Navier-Stokes equation as
@@
(4.13) $\mathrm{d}X_{t}=(A_{0}X_{t}+F(X_{t}))\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x.$
Existence and uniqueness of this equation has been discussed in [14, Example 2.9].
@@
The hemicontinuity has been proved in [14, Example 2.9].
@@
The coercivity has been proved [14, Example 2.9], with $\alpha=p$ for some $\delta_{1}>0$.
@@
By [14, Example 2.9], we get for all $u,v\in V$, that there exists $C>0$, such that for any $\varepsilon>0$, there exists $C(\varepsilon)>0$ with
@@
Hypothesis (D) holds by [14, Example 2.9].
@@
Note that by Korn’s inequality [58, Theorem 1.10 (p. 196)] and the previous inequality, for some $C>0$, which might change from line to line,
@@
see [14, Example 2.9]. Therefore, Hypothesis (E) is satisfied by Lemma 4.2, noting that $\langle F(u),u\rangle=0$ for $u\in V$.
Now, Theorem 2.5 follows from Theorem 2.1.
@@
5.1. Main estimates
We start with defining some constants.
Definition 5.1.
Assume that Hypotheses (A)(B)(C)(D), and (F) are valid. For any $\lambda_{i}\in(0,1)$, $i=1,2,3,4$, and $\lambda_{0}\in(0,1)$ with $\sum_{i=0}^{4}\lambda_{i}=1$, let
(5.1) $\begin{split}&c_{1}:=\frac{\beta+2}{2}\left(\frac{\beta}{(\alpha+\beta)\lambda_{1}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{\beta}{\alpha}}\frac{\alpha}{\alpha+\beta}\left(\left(1+18(\beta+2)\right)\|B\|^{2}_{2}+2\beta\|B\|_{L(H)}^{2}\right)^{\frac{\alpha+\beta}{\alpha}},\\ &c_{2}:=\left(\frac{2(\alpha-2)}{(\beta+2)(\alpha+\beta)\lambda_{2}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{2+\beta}{\alpha-2}}\frac{\alpha-2}{\alpha+\beta}(\gamma\|B\|_{2}^{2})^{\frac{\alpha+\beta}{\alpha-2}},\\ &c_{3}:=\frac{\beta+2}{2}\left(\frac{\beta}{(\alpha+\beta)\lambda_{1}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{\beta}{\alpha}}\frac{\alpha}{\alpha+\beta}(18(\beta+2)\mathfrak{g}_{0})^{\frac{\alpha+\beta}{\alpha}},\\ &c_{4}:=\left(\frac{2\beta}{(\beta+2)(\alpha+\beta)\lambda_{4}c_{0}^{\alpha}\delta_{1}}\right)^{\frac{\beta}{\alpha}}\frac{\alpha}{\alpha+\beta}(C_{\beta}\mathfrak{g}_{0})^{\frac{\alpha+\beta}{\alpha}},\\ &c_{5}:=C_{\beta}\mathfrak{g}_{\beta},\end{split}$
if $\alpha>2$, where
$C_{\beta}:=((2^{\beta-3}\vee 1)\beta(\beta+2))\vee((2^{\beta-1}\vee 1)(\beta+2)),\quad\beta>0,$
and $C_{0}:=\frac{1}{2}$ are as in Lemma A.1. Define $c_{1},c_{2},c_{4}$, and $c_{5}$ similarly if $\alpha=2$, and $c_{2}:=0$. Set
$\mathfrak{c}:=\sum_{i=1}^{5}c_{i}.$
Remark 5.2.
Note that $\mathfrak{c}=0$ if and only if $B\equiv G\equiv 0$, that is, we are considering the deterministic equation without noise. In this case, in the proof of Proposition 5.3 below, we can choose $\lambda_{0}=1$ and $\lambda_{i}=0$, $i=1,2,3,4$. In the pure Gaussian case, that is, $G\equiv 0$ and $B\not=0$, we may set $\lambda_{3}=\lambda_{4}=c_{3}=c_{4}=c_{5}=0$. Analogously, in the pure jump case $B\equiv 0$ and $G\not=0$, we may set $\lambda_{1}=\lambda_{2}=c_{1}=c_{2}=0$.
@@ 5.1. Main estimates
Assume that Hypotheses (A)(B)(C)(D), and (F) are valid. Let $T>0$ be fixed. Let $\alpha\geq 2,\beta\geq 0$ be as in Hypothesis (C). Then for any initial datum $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, and every $0\leq s\leq t\leq T$, the solution $(X_{t})_{t\in[0,T]}$ to (2.1) satisfies
(5.2) $\begin{split}&\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t}\|X_{r}\|^{\beta+2}_{H}\right]+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t}e^{\gamma(r-t)}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\\ \leq&\mathbb{E}\left[\|X_{s}\|_{H}^{\beta+2}e^{\gamma(s-t)}\right]+\frac{\mathfrak{c}}{\gamma}\left(1-e^{\gamma(s-t)}\right),\end{split}$
for some $\gamma\in(0,\delta_{2})$, where $\delta_{1}>0$ is as in Hypothesis (B), and $\delta_{2}>0$ is as in Hypothesis (C), and where $\lambda_{0}\in(0,1]$ and $\mathfrak{c}\geq 0$ are as in Definition 5.1, where $\lambda_{0}=1$ and $\mathfrak{c}=0$ if and only if $B\equiv G\equiv 0$. In particular, $\mathfrak{c}$ does not depend on $T$.
@@
In order to rigorously apply Itô’s formula, we construct approximating solutions to (2.1) via projections to finite dimensional subspaces using the so-called Faedo-Galerkin’s method.
In the sequel, we define the notation of the aforementioned projections. Let $\{e_{1},e_{2},\ldots,\}\subset V$ be an orthonormal basis for $H$ such that $\operatorname{span}\{e_{1},e_{2},\ldots,\}$ is dense in $V$. For each $n\in\mathbb{N}$, we denote $H_{n}:=\operatorname{span}\{e_{1},\ldots,e_{n}\}$ and set $Q_{n}:V^{\ast}\to H_{n}$ be the linear projection operator defined by
@@
We then denote for each $t\geq 0$
$W^{(n)}_{t}:=\sum_{j=1}^{n}\langle W_{t},e_{j}\rangle e_{j}=Q_{n}W_{t}.$
@@
(5.4) $\begin{split}\mathrm{d}X^{(n),x_{n}}_{t}&=Q_{n}A(X_{t}^{(n),x_{n}})\,\mathrm{d}t+Q_{n}B\,\mathrm{d}W_{t}+\int_{D^{c}}Q_{n}G(w)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}w),\quad t\geq 0,\\ X_{0}^{(n),x_{n}}&=Q_{n}x=:x_{n}.\end{split}$
@@
By (5.5) we have $\|X^{(n),x_{n}}_{t}\|_{V}\leq R$ for $0\leq t<\tau^{(n),x_{n}}_{R}$ and hence by Fact 3.1 we obtain $\|X^{(n),x_{n}}_{t}\|_{H}\leq c^{-1}_{0}R$ for $0\leq t<\tau^{(n),x_{n}}_{R}$. Similarly as in the proof of [14, Lemma 4.2] or in the proof of [69, Theorem 2.8], we have
@@
For the sake of readability, let us abbreviate $\tau_{R}:=\tau^{(n),x_{n}}_{R}$, and $X^{(n)}:=X^{(n),x_{n}}$. By Itô’s formula [38, 62, 39], or [74, Section 2.8], see also [14, proof of Lemma 4.2], it follows that $\mathbb{P}$-a.s.

### v4

@@ 4.3. Stochastic incompressible 2D Navier-Stokes equations
Let us prove Theorem 2.5. The deterministic incompressible 2D Navier-Stokes equations can be formulated as
@@
where $u:[0,T]\times{\mathcal{O}}\to\mathbb{R}^{2}$ denotes the velocity field of an incompressible Newtonian fluid with viscosity $\nu>0$, $p:[0,T]\times{\mathcal{O}}\to\mathbb{R}$ denotes the pressure, and $f:[0,T]\times{\mathcal{O}}\to\mathbb{R}^{2}$ denotes an external force, where ${\mathcal{O}}\subset\mathbb{R}^{2}$ is a domain with finite width with sufficiently smooth boundary, see [94, 1]. We employ no-slip (that is, Dirichlet) boundary conditions. Define
$V:=\left\{v\in W^{1,2}_{0}({\mathcal{O}};\mathbb{R}^{2})\;\colon\;\nabla\cdot v=0\text{ a.e. in }{\mathcal{O}}\right\}.$
@@
Denote the extensions $A_{0}:V\to V^{\ast}$ and $F:V\times V\to V^{\ast}$ by the same respective symbols. Note that these extensions can be defined by duality for Lipschitz boundary. Consider the stochastic incompressible 2D Navier-Stokes equations with viscosity $\nu>0$ driven by degenerate additive Wiener noise
(4.12) $\mathrm{d}X_{t}=(A_{0}X_{t}+F(X_{t}))\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}=x,$
where $B\in L_{2}(U,V)$. Existence and uniqueness of solutions to (2.9) have been discussed in [16, Example 2.6]. Let us verify our Hypotheses in the Gelfand triple
@@
Let us recollect Ladyzhenskaya’s inequality, see [64, 94],
$\langle F(u,v),u\rangle\leq 2\|u\|_{H}\|u\|_{V}\|v\|_{V},\quad u,v\in V.$
@@
The hemicontinuity has been proved in [16, Example 2.6].
@@
The coercivity has been proved [16, Example 2.6], with $\alpha=2$ and $\delta_{1}=2\nu$.
@@
Noting that by integration by parts and the incompressibility condition, $\langle F(u,v),v\rangle=0$ for any $u,v\in V$, we get by Ladyzhenskaya’s and Young’s inequalities, respectively, for all $u,v\in V$,
@@
$=$ $2\langle F(u,u-v),u-v\rangle+2\langle F(u-v,v),u-v\rangle$
$=$ $2\langle F(u-v,v),u-v\rangle$
$\leq$ $4\|u-v\|_{H}\|u-v\|_{V}\|v\|_{V}$
$\leq$ $\nu\|u-v\|_{V}^{2}+\frac{4}{\nu}\|v\|_{V}^{2}\|u-v\|_{H}^{2}.$
@@
$2\langle A_{0}u+F(u)-A_{0}v-F(v),u-v\rangle\leq-\nu\|u-v\|^{2}_{V}+\frac{4}{\nu}\|v\|_{V}^{2}\|u-v\|_{H}^{2}.$
We get that Hypothesis (B) holds with $\alpha=2$, $\beta=0$, $\delta_{2}=\nu c_{0}^{2}$ and $\rho(v)=\frac{4}{\nu}\|v\|_{V}^{2}$ and $C_{2}=\frac{4}{\nu}$.
@@
Hypothesis (D) holds by [16, Example 2.6].
@@
Then, Theorem 2.5 follows directly from Theorem 2.1 and Theorem 6.2.
Remark 4.4.
The above example includes the case of the deterministic 2D Navier-Stokes equations. We obtain that the unique invariant measure is a Dirac measure supported in $0\in H$, being the unique stationary solution $u_{\infty}$ to
$A_{0}u_{\infty}+F(u_{\infty})=0$
with Dirichlet boundary conditions. By taking $\langle\varphi,\cdot\rangle$, $\varphi\in H$ and $\|\cdot\|^{2}_{H}$ as test-functions for the semigroup, the mixing in $2$-Wasserstein distance yields strong convergence in $H$ to zero for the solutions to the deterministic 2D Navier-Stokes equations, as $t\to\infty$ (noting that, in Hilbert spaces, weak convergence, together with convergence of the norms, yields strong convergence). We recover the deterministic stability and extinction results in [94, Chapter 10].
@@ 4.4. Stochastic shear thickening incompressible power-law fluids
Let us prove Theorem 2.6. Consider the velocity field of a viscous and incompressible non-Newtonian fluid perturbed by Wiener noise with Dirichlet boundary conditions on a sufficiently smooth domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, with outward unit normal $n$ on $\partial{\mathcal{O}}$, $d\geq 2$. Let $p\geq 2$, $\nu>0$, and assume that $p\geq 1+\frac{d}{2}$. The case $p>2$ is called shear thickening. For $u:{\mathcal{O}}\to\mathbb{R}^{d}$, define
@@
and define the convection term $F:(W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V)\times(W^{2,p}({\mathcal{O}};\mathbb{R}^{d})\cap V)\to H$ as before in the case of the 2D Navier-Stokes equations as
@@
(4.13) $\mathrm{d}X_{t}=(A_{0}X_{t}+F(X_{t}))\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}=x.$
Existence and uniqueness of this equation has been discussed in [16, Example 2.9].
@@
The hemicontinuity has been proved in [16, Example 2.9].
@@
The coercivity has been proved [16, Example 2.9], with $\alpha=p$ for some $\delta_{1}>0$.
@@
By [16, Example 2.9], we get for all $u,v\in V$, that there exists $C>0$, such that for any $\varepsilon>0$, there exists $C(\varepsilon)>0$ with
@@
Hypothesis (D) holds by [16, Example 2.9].
@@
Note that by Korn’s inequality [73, Theorem 1.10 (p. 196)] and the previous inequality, for some $C>0$, which might change from line to line,
@@
see [16, Example 2.9]. Therefore, Hypothesis (E) is satisfied by Lemma 4.2, noting that $\langle F(u),u\rangle=0$ for $u\in V$.
For Hypothesis (H), we need that, $\alpha\geq 2$, $\beta\geq 0$ and that $\alpha-\beta-2\geq 0$ which is equivalent to $\geq d, p2p\geq 2$, $2-2p+d\geq 0$, and $(2p-2)(2p-d)\geq 2p$ in this situation. In particular, $d\geq 2$ and $p\in[\frac{d}{2},1+\frac{d}{2}]$. Hence, as $p\geq 1+\frac{d}{2}$, we remain only with the case that $p=\alpha=1+\frac{d}{2}$, $\beta=0$. Note that $(2p-2)(2p-d)\geq 2p$ is automatically satisfied for $p=1+\frac{d}{2}$ and $d\geq 2$. Now, Theorem 2.6 follows from Theorem 2.1 and Theorem 6.2.
@@
5.1. Main estimate
@@
Assume that Hypotheses (A)(B)(C)(D)(F). Let $T>0$ be fixed. Let $\alpha\geq 2$, $\alpha-2\geq\beta\geq 0$ be as in Hypothesis (C). If $\alpha=\beta+2$, we assume additionally that
$\|B\|_{L_{2}(U,H)}^{2}\leq\lambda_{3}\frac{1}{\alpha}\delta_{1}c_{0}^{\alpha},$
where $\lambda_{i}$, $i=0,1,2,3$, are as in Hypothesis (H), and $\delta_{1}>0$ is as in Hypothesis (B). Then for any initial datum $X_{0}=x$ such that $\exp(\|x\|^{\beta+2}_{H})\in L^{1}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, and every $0\leq s\leq t\leq T$, the solution $(X_{t})_{t\in[0,T]}$ to (2.1) satisfies
(5.1) $\begin{split}&\mathbb{E}\left[\exp\left(\|X_{t}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\right)\right]\\ \leq&\mathbb{E}\left[\exp\left(\|X_{s}\|_{H}^{\beta+2}+(t-s)(c_{1}+c_{2}+c_{3})\right)\right],\end{split}$
for every $0\leq s\leq t\leq T$ in the locally monotone case, and similarly with $\mathbb{E}$ replaced by $\tilde{\mathbb{E}}$ in the fully locally monotone case. Here, $c_{i}\geq 0$, $i=1,2,3$, are as in Hypothesis (H). Furthermore,
(5.2) $\begin{split}&\mathbb{E}\left[\exp\left(\|X_{t}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{0}^{t}\|X_{r}\|^{\beta}_{H}\|X_{r}\|^{\alpha}_{V}\,\mathrm{d}r\right)\right]\\ \leq&\mathbb{E}\left[\exp\left(\|x\|_{H}^{\beta+2}+t(c_{1}+c_{2}+c_{3})\right)\right].\end{split}$
for every $0\leq t\leq T$ in the locally monotone case, and similarly with $\mathbb{E}$ replaced by $\tilde{\mathbb{E}}$ in the fully locally monotone case.
@@
In order to rigorously apply Itô’s formula, we construct approximating solutions to (2.1) via projections to finite dimensional subspaces using the Faedo-Galerkin’s method.
In the sequel, we define the notation of the aforementioned projections. Let $\{e_{1},e_{2},\ldots,\}\subset V$ be a complete orthonormal basis for $H$ such that $\operatorname{span}\{e_{1},e_{2},\ldots,\}$ is dense in $V$. For each $n\in\mathbb{N}$, we denote $H_{n}:=\operatorname{span}\{e_{1},\ldots,e_{n}\}$ and set $Q_{n}:V^{\ast}\to H_{n}$ be the linear projection operator defined by
@@
Let $\{g_{1},g_{2},\ldots,\}\subset U$ be a complete orthonormal basis for $U$. We then denote for each $t\geq 0$
$W^{(n)}_{t}:=\tilde{Q}_{n}W_{t}:=\sum_{j=1}^{n}\langle W_{t},g_{j}\rangle g_{j}.$
@@
(5.4) $\begin{split}\mathrm{d}X^{(n),x_{n}}_{t}&=Q_{n}A(X_{t}^{(n),x_{n}})\,\mathrm{d}t+Q_{n}B\,\mathrm{d}W^{(n)}_{t},\quad t\geq 0,\\ X_{0}^{(n),x_{n}}&=Q_{n}x=:x_{n}.\end{split}$
@@
By (5.5) we have $\|X^{(n),x_{n}}_{t}\|_{V}\leq R$ for $0\leq t<\tau^{(n),x_{n}}_{R}$ and hence by Fact 3.1 we obtain $\|X^{(n),x_{n}}_{t}\|_{H}\leq c^{-1}_{0}R$ for $0\leq t<\tau^{(n),x_{n}}_{R}$. Similarly as in the proof of [83, Theorem 2.8], we have
@@
For the sake of readability, let us abbreviate $\tau_{R}:=\tau^{(n),x_{n}}_{R}$, and $X^{(n)}:=X^{(n),x_{n}}$. By Itô’s formula [50, 77, 51], or [90, Section 2.8], it follows that $\mathbb{P}$-a.s.